OpenStudy (anonymous):

how do you solve \sqrt{2m-6}=\sqrt{3m-14}

OpenStudy (anonymous):

it's really simple if you just square both sides

OpenStudy (anonymous):

m=8

OpenStudy (anonymous):

i did but i got it wrong

OpenStudy (anonymous):

is it 2m-6=3m-14 2m+8=3m 8=m

OpenStudy (anonymous):

\[(\sqrt{2m-6})^{2}=(\sqrt{3m-14})^{2}\]

OpenStudy (anonymous):

2m-6=3m-14 -m=-8 m=8

OpenStudy (anonymous):

yeah thats what i put but my friend said it was wrong and how did i get that

OpenStudy (anonymous):

what the reason??

OpenStudy (anonymous):

just to show how i got that like step by step

OpenStudy (radar):

Step 1. Square both sides getting: 2m-6=3m-14 Step 2. add 6 and -3m to both sides of the equation getting -1m=-8 Step 3. divide both sides by -1 getting m=8. asrofie laid it out for you step by step.

OpenStudy (anonymous):

thank you

OpenStudy (radar):

What say naynay?

OpenStudy (anonymous):

thank you i get it now

OpenStudy (radar):

Great, good luck with your studies.